OpenStudy (anonymous):

Is the summation of 1/sqrt(n^3 + 1) , from 1 to infinity, convergent?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh this was not a yes or no question was it? it is yes converges

OpenStudy (anonymous):

quick eyeball check says \[\sqrt{n^3}=n^{\frac{3}{2}}\] and since \[\frac{3}{2}>1\] you win

OpenStudy (anonymous):

you want a test? integral test might do it but i think even comparison test works. compare to what i wrote

OpenStudy (anonymous):

Just to clarify. Should I compare \[1/\sqrt{n^3 + 1}\] with \[1/\sqrt{n^3}\] ??

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

what is a 1 between friends?

OpenStudy (anonymous):

So I'm comparing it with something divergent (by P-Series)? o.O

OpenStudy (anonymous):

how pedantic do you want to be? you could use the integral test, which would involve showing \[\int_1^{\infty}\frac{1}{\sqrt{x^3+1}}dx\] was finite,but this integral is a drad

OpenStudy (anonymous):

oh no p series converges if p> 1 right?

OpenStudy (anonymous):

\[\sum\frac{1}{n^p}\] is finite if p > 1

OpenStudy (anonymous):

Ok. I got mixed up with 3/2 and 2/3. Thanks. xD